Work the following problem using the algorithm we covered in class. Show your work by listing all the steps, Use Example 13.8 on page 706.

For the following reaction at 600 °C: 2SO2 (g) + O2 (g) ⇌ 2SO3(g) Kc = 4.32

What are the equilibrium concentrations of all species in a mixture that was prepared with [SO3] = 0.500 M, [SO2] = 0 M, and [O2] = 0.350 M?
Answer #1

2502(g) + O2(g) = 2503(9)

Using the ICE table

[SO2] [O2] [SO3]
Initial 0 0.350 0.500
Change +2x +x -2x
Final 2x 0.350+x 0.500-2x

Kc = [Products [Reactants S0312 (SO2)2 02]

(0.500-2.) (2x)2(0.350+) = 4.32

Solving the equation we get the value of x

I=0.104

[SO3] = 0.500 - 2(0.104) = 0.292

[O2] = 0.350 + 0.104 = 0.454

[SO2] = 2(0.104) = 0.208

Note - Post any doubts/queries in comments section.